OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 112 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t^2 , where h represents height in feet and r represents the rate in feet per secound

OpenStudy (anonymous):

when the ball hits the ground h = 0 0 = 112 t - 16t^2 t(112 - 16t) = 0 16t = 112 t = 7 seconds

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

which part don't you understand?

OpenStudy (anonymous):

The whole thing

OpenStudy (anonymous):

h = 0 because the ball is on the ground 112 is the value for r - the initial speed of the ball so these are plugged into the given formula t represents the time which we need to find so we rearrange the formula to make t the subject t(112 - 16t) = 0 either t = 0 or 112-16t = 0 which gives t = 7 seconds the t = 0 comes from when batter first hits the ball 7 seconds is the time that the ball is in the air until it hits the ground again

OpenStudy (anonymous):

Oh I get it now do u know what a box mean

OpenStudy (anonymous):

box?

OpenStudy (anonymous):

best answer box?

OpenStudy (anonymous):

No its a set with a box in it